OpenStudy (anonymous):

5-3i divided by 1-2i

OpenStudy (anonymous):

\[\frac{5-3i}{1-2i}\cdot\frac{1+2i}{1+2i}=\frac{5-3i+10i-6i^2}{1-4i^2}=\frac{11+7i}{5}\]

OpenStudy (anonymous):

thank youu!!

OpenStudy (anonymous):

do you think you can help me with more equation